# A zero dimensional study of Ohmically heated negative triangularity tokamaks

### v1

@@ Conclusions
In this work, we used 0D power balance to analyze the amount of heating power needed for a NT tokamak to produce fusion power and achieve large fusion gain. In this framework, we focused on NT plasmas with Ohmic heating only, which is the lowest possible amount of heating (as it is necessarily generated by the plasma current). While reducing the heating power did not enable more fusion power, it can allow a machine to achieve higher fusion gain $Q$ and minimize recirculating power. We made several comparisons with four machines characterized by different size, magnetic field and fusion gain: MANTA, SPARC, ITER and DEMO. We observed that NT ohmic operations are an attractive option for MANTA, SPARC and DEMO, while they are detrimental for ITER. These comparisons indicated that an Ohmic NT scenario is most attractive for design points with high fusion gain ($Q\gtrsim 10)$, especially when achieved using a strong magnetic field (as opposed to large size).
Acknowledgments

### v2

@@ Conclusions
Since NT tokamaks are not constrained to exceed the L-H heating power threshold, they can improve their energy confinement time by reducing their external heating. While reducing the heating power does not enable more fusion power, it can enable NT tokamaks to achieve higher fusion gain $Q$ and minimize recirculating power. We observed that NT Ohmic operations are an attractive option for MANTA, SPARC and DEMO, while they are detrimental for ITER. These comparisons along with analytic calculations indicated that an Ohmic NT scenario is most attractive for design points with high fusion gain ($Q\gtrsim 10)$, especially when achieved using a strong magnetic field (as opposed to large size).
I Acknowledgments
@@
II Appendix A
In this appendix we give explicit expressions for all the terms on the right side of equation (1). All the following formulas give power densities expressed in units of $keV\,s^{-1}\,m^{-3}$. The $\alpha$ power density is given by [29]
where $E_{\alpha}=3500\;keV$ is the energy of the $\alpha$ particles produced by deuterium-tritium fusion reactions, $n$ is expressed in $m^{-3}$, $C_{0}=6.4341\times 10^{-20}$, $\xi=C_{1}/(T^{1/3})$, $\zeta=1-(C_{2}T+C_{4}T^{2}+C_{6}T^{6})/(1+C_{3}T+C_{5}T^{2}+C_{7}T^{3})$, $C_{1}=6.661\,keV^{1/3}$, $C_{2}=1.5136\times 10^{-2}\,keV^{-1}$, $C_{3}=7.5189\times 10^{-2}\,keV^{-1}$, $C_{4}=4.6064\times 10^{-3}\,keV^{-2}$, $C_{5}=1.35\times 10^{-2}\,keV^{-2}$, $C_{6}=-1.0675\times 10^{-4}\,keV^{-3}$, $C_{7}=1.366\times 10^{-5}\,keV^{-3}$. The Ohmic power density is
where $\eta$ is the resistivity, $j$ is the current density, $\eta_{S}=1.64\times 10^{8}Z_{eff}\,m\,s^{-1}\,A^{-2}\,keV^{2.5}$ is the Spitzer resistivity coefficient, $Z_{eff}$ is the plasma effective charge, $\epsilon=a/R$ is the inverse aspect ratio, $I_{p}$ is the plasma current in $A$, $\mathcal{F}(\kappa_{95},\delta_{95},\epsilon)=4.1\times 10^{6}(1+1.2(\kappa_{95}-1)+0.56(\kappa_{95}-1)^{2})(1+0.09\delta_{95}+0.16\delta_{95}^{2})\frac{1+0.45\delta_{95}\epsilon}{1-0.74\epsilon}$ is a function that depends only on geometry and is used to correctly express the current density in terms of $I_{p}$ [30] for shaped plasmas, $\kappa_{0}$ is the elongation on axis, and $\kappa_{95}$ and $\delta_{95}$ are the elongation and triangularity at the flux surface enclosing 95% of the poloidal flux. The radiated power density is
where $C_{B}=3.3\times 10^{-21}\;m^{3}\,sec^{-1}\,keV^{0.5}$. Here, we are considering only Bremsstrahlung radiation, as it is typically the dominant contribution to the power lost by radiation. The power density lost through transport is extraordinarily difficult to predict from first-principles as it is dominated by turbulence. For this reason, we will use empirical scaling laws based on experimental results for the energy confinement time $\tau_{E}$ [31]. This enables us to write the loss term as
where the factor $3$ comes from the assumption that electrons and ions have the same temperature and density. Finally, the external power density $\langle p_{ext}\rangle$ is assumed to be a constant specified value, regardless of the plasma properties.
III Appendix B
For the Ohmic energy confinement time, we use the standard Neo-Alcator scaling [31]
or the ITER-89P L-mode scaling [33]
where $I_{p}$ is the plasma current in $MA$, $n_{19}$ and $n_{20}$ are the electron density in $10^{19}m^{-3}$ and $10^{20}m^{-3}$ respectively, $B$ is the magnetic field in $T$, $M$ is the average ion mass in atomic mass units, $R$ and $a$ are the major and minor radii in $m$, $\kappa_{a}$ is the elongation at the separatrix, and $P=P_{\alpha}+P_{\Omega}+P_{ext}$ is the total heating power in $MW$. It is important to note that the negative exponent in the heating power factor means that less external heating improves confinement. Like $H_{NA}$, the confinement enhancement factors $H_{98}$ and $H_{89}$ are scalars representing how much better or worse confinement is with respect to the nominal scaling law (i.e. $H_{98}=1$ or $H_{89}=1$). Since no existing scaling law incorporates NT with sufficiently broad applicability, one is forced to account for NT by choosing representative confinement enhancement factors. These three confinement factors represent the biggest uncertainty in predicting how a NT plasma will behave. A large number of NT discharges produced on TCV (figure 4(a)) and DIII-D [9, 18] indicate that the H-mode scaling of equation (12) with $H_{98}=1$ is a good fit for the experimental performance of heated NT plasmas. However, we will also consider a range of possiblities by using the ITER-89P scaling with different values of $H_{89}$. Regarding the Ohmic confinement time, a preliminary analysis of the TCV NT database suggest that $H_{NA}=2$ can be easily achieved (figure 4(b)) and thus this is our choice for our analysis.
Figure 4: (a) ITER98 and (b) Neo-Alcator confinement enhancement factors as functions of normalized $\beta$ and $q_{95}$ from sampled NT TCV shots with external heating (a) and Ohmic heating only (b).